OpenStudy (anonymous):

The Perimeter of a Rectangular house is 122 ft. The Width is 5ft less than the length. Find the length and the width. For an equation, I wrote: (5-w)+(5-w)+l+l=122 Now im having problems solving it. Im not sure what Im doing wrong. PLEASE HELP!!!

myininaya (myininaya):

l is length and w=l-5 where w is width

myininaya (myininaya):

2l+2(l-5)=122

OpenStudy (anonymous):

Oh, ok! I think I get it now! That's the first part of my equation.

myininaya (myininaya):

let me know what you get for l and w i will let you if you are right or not

OpenStudy (anonymous):

Ok thanks! Hold on . . .

OpenStudy (anonymous):

I think i got lost! I then got: 4l (4l)=122 Now, Im confused!

myininaya (myininaya):

ok we have 2l+2l-10=122 4l-10=122 4l=122+10=132 4l=132 l=132/4=33 so w=l-5=33-5=28

OpenStudy (anonymous):

ok wait! How did you get 10?

myininaya (myininaya):

2l+2(l-5)=122 i distribute the two

OpenStudy (anonymous):

LOL! I totally got it now! Im laughing cuz I forgot to distribute on my math test last week! OK! It all makes sense now! Thanks!

myininaya (myininaya):

cool